# Review-based Recommender Systems: A Survey of Approaches, Challenges and Future Perspectives

### v2

@@ 5.2. Deep Learning-based Approaches
Lastly, Asymmetrical Hierarchical Networks (AHN) (Dong et al., 2020) and Reliable Recommendation with Review-level Explanations (RRRE) (Lyu et al., 2021) focus on the granularity and reliability of reviews respectively. AHN addresses the varying scope of user and item reviews by applying differently focused attention modules, enhancing the relevancy of content extracted for recommendations.Specifically, the model utilizes bi-directional LSTM networks to process the sequential data within reviews, allowing for the effective capture and integration of contextual and temporal relationships between sentences. Conversely, RRRE prioritizes the authenticity of reviews, using mechanisms to assess and filter content based on reliability, which is crucial for maintaining the integrity of recommendation outputs. Overall, these RNN-based models improve recommendation accuracy and user satisfaction by processing sequential review data. They vary in their specific applications—some enhance interpretability and trust (TARMF and RRRE), while others focus on adapting to changing user conditions (DER and RRGM). The use of RNNs alongside technologies like attention mechanisms and adversarial networks indicates the potential for ongoing advancements in the field, aiming to align recommender systems more closely with user expectations and preferences.
Graph Neural Network-based Approaches: Graph Neural Networks (GNNs) have significantly advanced in dealing with complex user-item interactions in recommendation systems. By exploiting the structural connectivity of data, GNNs excel in capturing high-order relationships. This is particularly effective for addressing data sparsity, cold start problems, and improving the interpretability of recommendations through detailed insights from user and item interactions.
Tripartite Graph Ranking (TrinRank) (He et al., 2015) evolves beyond traditional collaborative filtering by modeling ternary relationships in a heterogeneous tripartite graph (user-item-aspect). The model focuses on vertex ranking in the tripartite graphs, tailored for personalized recommendation which provides explainability and transparency by modeling aspects in reviews. In contrast, Review Meet Graphs (RMG) (Wu et al., 2019c) utilizes a multi-view framework that merges review texts with user-item graphs within a GNN architecture. By doing so, RMG captures detailed user-item interactions more adeptly, highlighting the advantage of integrating diverse data views to boost the effectiveness of recommender systems.
@@
Contrastive Learning-based Approaches: The Contrastive Learning-based Approach (Khosla et al., 2020; Xia et al., 2022) refines recommender systems by its unique power of data representation learning technique - modeling the data representation by minimizing the distance between similar data points while maximizing the distance between similar data points. This approach has been successful in tasks requiring detailed data understanding, and making recommendations based on user reviews.
The Review-aware Graph Contrastive Learning (RGCL) framework leads this approach by enhancing user and item embeddings through the integration of review data. It utilizes contrastive learning techniques to improve node embedding and interaction modeling, thus addressing data sparsity and more accurately capturing user preferences. Similarly, the Review-based Multi-intention Contrastive Learning (RMCL) (Yang et al., 2023) method introduces learning intention representations from reviews based on a mixed Gaussian distribution hypothesis. It uses a multi-intention contrastive learning strategy to link user and item reviews, fostering a deeper connection between them.
@@
LLMs-based Approaches: Large Language Models (LLMs) excel at processing complex text, which enhances the personalization and clarity of recommender systems (Laskar et al., 2023a, b). A significant advancement in this area has been achieved by integrating LLMs, specifically GPT-2, into the recommendation process (Li et al., 2023b). By utilizing a prompt tuning technique, the model dynamically extracts key aspects reflecting user preferences and item properties directly from textual reviews. These extracted aspects are then used to inform and personalize the recommendation process, leading to a more tailored and insightful user experience.
This integration not only improves the relevance and quality of recommendations but also contributes to the transparency and explainability of the system, addressing a critical need in recommender systems research. Evaluated on three large-scale datasets, the model demonstrates superior performance over existing methods, showcasing its potential for practical application and scalability. This study demonstrates the transformative potential of LLMs in recommender systems, paving the way for further exploration and development in this exciting intersection of natural language processing and personalized recommendation.
@@
Category Models Year Architecture Features Datasets Evaluation Tasks
ConvMF(Kim et al., 2016) 2016 CNN, MF reviews amazon, yelp & beer RMSE accuracy & interpretability
@@
D-Attn(Seo et al., 2017) 2017 CNN, Attn reviews yelp & amazon MSE accuracy & interpretability
NARRE(Chen et al., 2018a) 2018 CNN, Attn reviews yelp & amazon MSE accuracy
SentiRec(Hyun et al., 2018) 2018 CNN reviews amazon MSE scalability & memory efficient
@@
TARMF(Lu et al., 2018) 2018 GRU, Attn ratings & reviews amazon & yelp MSE accuracy
AIN(Mei et al., 2018) 2018 Attn ratings & reviews yelp, food & frappe RMSE accuracy & interpretability
ANR(Chin et al., 2018) 2018 Attn reviews yelp & amazon MSE, MAE accuracy & interpretability
SDNet(Chen et al., 2018b) 2018 CNN, LSTM review amazon MSE accuracy
DL CARL(Wu et al., 2019a) 2019 CNN, Attn reviews amazon MSE accuracy
@@
DAML(Liu et al., 2019a) 2019 CNN, MF reviews amazon MAE accuracy & interpretability
AARM(Guan et al., 2019) 2019 Attn reviews yelp & amazon N & HR accuracy & interpretability
@@
SDRA(Da’u and Salim, 2019) 2019 LSTM reviews amazon & yelp MSE accuracy
@@
RSBM(Huang et al., 2020b) 2020 CNN, Attn reviews amazon & yelp MSE accuracy
@@
AODR(Da’u et al., 2020) 2020 CNN reviews amazon MSE, RMSE accuracy
SSG(Gao et al., 2020b) 2020 CNN review amazon & yelp MAE accuracy
NCTR(Liu et al., 2021a) 2021 CNN reviews & ratings amazon & movie MSE, MAE accuracy
@@
Acapnet(Yang et al., 2021) 2021 CNN reviews amazon MSE accuracy
MrRec(Liu et al., 2021d) 2021 CNN, Attn reviews amazon & goodreads MSE acc, diversity & cold start
SIFN(Zhang et al., 2021) 2021 Attn review amazon MSE accuracy
DRRNN(Xi et al., 2021) 2021 CNN, MLP reviews & ratings amazon RMSE data sparsity
U-ARM(Sun et al., 2021) 2021 DNN reviews amazon & yelp RMSE accuracy & interpretability
ERP(Wu et al., 2021) 2021 CNN reviews & ratings amazon & yelp MSE, RMSE accuracy
@@
RPR(Liu et al., 2022a) 2022 CNN reviews amazon & yelp MSE & MAE accuracy
@@
MCCL(Wei et al., 2024) 2024 GNN & CL reviews amazon & yelp NDCG, P accuracy
Reinforcement Learning-based Approaches: In Reinforcement Learning (RL), an agent learns to make decisions by interacting with an environment (Wiering and Van Otterlo, 2012). The agent receives rewards or penalties based on its actions, which guide it toward optimizing its behavior over time. In review-based recommender systems (RBRS), RL can be particularly effective in scenarios where user preferences are dynamic and the system must adapt to continuously evolving data. RL algorithms can optimize the recommendation process by learning from user interactions and feedback, such as reviews, to enhance the personalization and accuracy of the recommendations (Afsar et al., 2022).
@@
This research opens up areas for exploring reinforcement learning in review-based recommender systems. Future work could focus on developing more sophisticated RL algorithms to enhance the system’s resilience. Additionally, exploring the integration of RL with other machine learning techniques could lead to more robust and adaptable recommender systems that can handle increasingly complex user interactions and preferences.
Miscellaneous Approaches: The Sentiment Utility Logistic Model (SULM) (Bauman et al., 2017) utilizes sentiment analysis to identify key aspects of products that influence user preferences, enhancing the personalization of recommendations. For example, it might recommend a restaurant specifically for its standout seafood, effectively tailoring suggestions to user preferences. Demonstrated in various settings like restaurants and hotels, SULM’s aspect-focused recommendations have been well-received for improving user satisfaction. Building on this focus on user insights, the Adversarial Training for Review-based Recommendations (ATR) (Rafailidis and Crestani, 2019) model integrates user reviews with collaborative filtering. By using adversarial training, this model aligns deep representations of reviews with user and item features, aiming to improve the accuracy of recommendations.
@@
The Sentiment-aware Interactive Fusion Network (SIFN) (Zhang et al., 2021) addresses the precision of sentiment interpretation in recommendations by utilizing Bidirectional Encoder Representation From Transformer (BERT) (Devlin et al., 2018) for encoding and a lightweight sentiment learner for semantic analysis. This model enhances the accuracy of item suggestions by using explicit sentiment labels, although it faces challenges with sparse or vague reviews. Counterfactual Explainable Recommendation (CountER) (Tan et al., 2021) leverages counterfactual reasoning to provide clear and meaningful explanations for recommendations by identifying minimal changes in item aspects that could influence user decisions. CountER not only enhances the transparency of recommendations but also assesses the impact and clarity of its explanations, significantly outperforming other methods in explainability.
@@
The Lightweight Unbiased Multi-teacher Ensemble (LUME) (Xv et al., 2022) addresses computational complexity, biased recommendations, and poor generalization by implementing a multi-teacher ensemble with debiased knowledge distillation. This strategy aggregates insights from various pretrained models to form a streamlined, unbiased system that excels in generalization. The detailed summary of the deep learning-based approaches are presented in Table 3.
6. Datasets
Traditional rating-based recommender systems rely on the dataset that comes with user, item, and numerical rating information, e.g. MovieLens Dataset  [^4] https://grouplens.org/datasets/movielens/, IMDb Dataset [^5] https://developer.imdb.com/non-commercial-datasets/, and Goodreads Dataset [^6] https://www.goodreads.com/genres/dataset. In contrast, review-based recommender models are evaluated mainly with the datasets that contain user, item, and users’ textual review feedback, e.g. Amazon [^7] https://jmcauley.ucsd.edu/data/amazon/, Yelp [^8] https://www.yelp.com/dataset, TripAdvisor [^9] https://www.cs.virginia.edu/~hw5x/Data/LARA/TripAdvisor/, and Beer reviews  [^10] https://cseweb.ucsd.edu/~jmcauley/datasets.html#multi_aspect.

### v3

@@ 5.2.4. Recurrent Neural Networks (RNNs)-based Approaches:
Lastly, Asymmetrical Hierarchical Networks (AHN) (Dong et al., 2020) and Reliable Recommendation with Review-level Explanations (RRRE) (Lyu et al., 2021) focus on the granularity and reliability of reviews respectively. AHN addresses the varying scope of user and item reviews by applying differently focused attention modules, enhancing the relevancy of content extracted for recommendations. Specifically, the model utilizes bi-directional LSTM networks to process the sequential data within reviews, allowing for the effective capture and integration of contextual and temporal relationships between sentences. Conversely, RRRE prioritizes the authenticity of reviews, using mechanisms to assess and filter content based on reliability, which is crucial for maintaining the integrity of recommendation outputs. Overall, these RNN-based models improve recommendation accuracy and user satisfaction by processing sequential review data. They vary in their specific applications—some enhance interpretability and trust (TARMF and RRRE), while others focus on adapting to changing user conditions (DER and RRGM). The use of RNNs alongside technologies like attention and adversarial networks indicates the potential for ongoing advancements in the field, aiming to align recommender systems more closely with user expectations and preferences.
5.2.5. Graph Neural Network (GNN)-based Approaches:
GNNs can model the complex user-item interactions in recommendation systems. By exploiting the structural connectivity of data, GNNs excel in capturing high-order relationships. This is particularly effective for addressing data sparsity, cold start problems, and improving the interpretability of recommendations through detailed insights from user and item interactions. Tripartite Graph Ranking (TrinRank) (He et al., 2015) evolves beyond traditional collaborative filtering by modeling ternary relationships in a heterogeneous tripartite graph (user-item-aspect). The model focuses on vertex ranking in the tripartite graphs, tailored for personalized recommendation which provides explainability and transparency by modeling aspects in reviews. In contrast, Review Meet Graphs (RMG) (Wu et al., 2019c) utilizes a multi-view framework that merges review texts with user-item graphs within a GNN architecture. By doing so, RMG captures detailed user-item interactions more adeptly, highlighting the advantage of integrating diverse data views to boost the effectiveness of recommender systems.
@@ 5.2.5. Graph Neural Network (GNN)-based Approaches:
[…]
The Evolution of the Review-based Recommendation Between 2015 to 2024
Figure 5. The Evolution of the Review-based Recommendation Between 2015 to 2024
5.2.6. Contrastive Learning-based Approaches:
The Contrastive Learning-based Approach (Khosla et al., 2020; Xia et al., 2022) refines recommender systems by its unique power of data representation learning technique - modeling the data representation by minimizing the distance between similar data points while maximizing the distance between dissimilar data points. This approach has been successful in tasks requiring detailed data understanding, and making recommendations based on user reviews. The review-aware Graph Contrastive Learning (RGCL) framework leads this approach by enhancing user and item embeddings through the integration of review data. It utilizes contrastive learning techniques to improve node embedding and interaction modeling, thus addressing data sparsity and more accurately capturing user preferences. Similarly, the Review-based Multi-intention Contrastive Learning (RMCL) (Yang et al., 2023a) method introduces learning intention representations from reviews based on a mixed Gaussian distribution hypothesis. It uses a multi-intention contrastive learning strategy to link user and item reviews, fostering a deeper connection between them.
@@ 5.2.6. Contrastive Learning-based Approaches:
5.2.7. LLMs-based Approaches:
Large Language Models (LLMs) excel at processing complex text, which enhances the personalization and clarity of recommender systems (Laskar et al., 2023a, b). A significant advancement in this area has been achieved by integrating LLMs, specifically GPT-2, into the recommendation process (Li et al., 2023b). By utilizing a prompt tuning technique, the model dynamically extracts key aspects reflecting user preferences and item properties directly from textual reviews. These extracted aspects are then used to inform and personalize the recommendation process, leading to a more tailored and insightful user experience.
This integration not only improves the relevance and quality of recommendations but also contributes to the transparency and explainability of the system, addressing a critical need in recommender systems research. Evaluated on three large-scale datasets, the model demonstrates superior performance over existing methods, showcasing its potential for practical application and scalability. This study demonstrates the transformative potential of LLMs in recommender systems, paving the way for further exploration and development in this exciting intersection of natural language processing and personalized recommendation. The Sentiment-aware Interactive Fusion Network (SIFN) (Zhang et al., 2021) addresses the precision of sentiment interpretation in recommendations by utilizing Bidirectional Encoder Representation From Transformer (BERT) (Kenton and Toutanova, 2019) for encoding and a lightweight sentiment learner for semantic analysis. This model enhances the accuracy of item suggestions by using explicit sentiment labels, although it faces challenges with sparse or vague reviews.
@@ 5.2.7. LLMs-based Approaches:
Models Year Architecture Features Datasets Evaluation Tasks
TriRanK (He et al., 2015) 2015 GNN reviews amazon HR, NDCG transparency, explainability
ConvMF (Kim et al., 2016) 2016 CNN, MF reviews amazon, MovieLens RMSE accuracy
@@
RDSA (Preethi et al., 2017) 2017 RNN reviews amazon F1 accuracy
D-Attn (Seo et al., 2017) 2017 CNN, Attn reviews yelp & amazon MSE accuracy
SULM (Bauman et al., 2017) 2017 SULM reviews yelp Precision@K accuracy, interpretability
NARRE (Chen et al., 2018a) 2018 CNN, Attn reviews yelp & amazon MSE accuracy, explainability
SentiRec (Hyun et al., 2018) 2018 CNN reviews amazon MSE scalability, memory efficient
@@
TARMF (Lu et al., 2018) 2018 GRU, Attn ratings & reviews amazon & yelp MSE transparency, explainability
AIN (Mei et al., 2018) 2018 Attn ratings & reviews yelp, food & frappe RMSE accuracy, interpretability
ANR (Chin et al., 2018) 2018 CNN reviews yelp & amazon MSE, MAE accuracy, interpretability
PARL (Wu et al., 2018) 2018 CNN reviews beer & amazon MSE accuracy
MCPN (Tay et al., 2018) 2018 Attn reviews yelp & amazon MSE, MAE accuracy, interpretability
SDNet (Chen et al., 2018b) 2018 CNN, LSTM reviews amazon MSE accuracy, time complexity
CARL (Wu et al., 2019a) 2019 CNN, Attn reviews amazon MSE accuracy, interpretability
@@
DAML (Liu et al., 2019a) 2019 CNN, MF reviews amazon MAE accuracy, interpretability
AARM (Guan et al., 2019) 2019 Attn reviews yelp & amazon N & HR accuracy, interpretability
RSBM (Huang et al., 2020b) 2019 CNN, Attn reviews amazon & yelp MSE accuracy, scalability
AODR (Da’u et al., 2020) 2019 CNN reviews amazon MSE, RMSE accuracy
@@
SDRA (Da’u and Salim, 2019b) 2019 LSTM reviews amazon & yelp MSE accuracy
@@
ACMLM (Ni et al., 2019) 2019 GNN reviews amazon, yelp RMSE accuracy
RRGM (Xia et al., 2019) 2019 GRU, Attn reviews amazon, yelp MSE accuracy, interpretability
RMG (Wu et al., 2019c) 2019 GNN ratings & reviews amazon, yelp RMSE accuracy
DER (Chen et al., 2019) 2019 CNN, GRU ratings & reviews amazon, yelp RMSE accuracy, explainability
CRARec (Shoja and Tabrizi, 2019) 2019 LDA, MLP ratings & reviews amazon MSE accuracy
@@
HGNR (Liu et al., 2020) 2020 GNN reviews yelp, L., eps MSE, MAE accuracy
SSG (Gao et al., 2020b) 2020 CNN reviews amazon & yelp HR, NDCG accuracy
SARecS (Hung, 2020) 2020 CNN-LSTM ratings & reviews amazon RMSE accuracy
AHN (Dong et al., 2020) 2020 Bi-LSTM, Attn ratings & reviews amazon MAE accuracy, interpretability
EHRecSA (Ray et al., 2020) 2020 BERT reviews amazon & yelp P, R, F accuracy
MSAR (Peng et al., 2020) 2020 CNN, Attn ratings & reviews amazon MSE accuracy
NCTR (Liu et al., 2021a) 2020 CNN ratings & reviews amazon & movie MSE, MAE accuracy
@@
Acapnet (Yang et al., 2021) 2021 CNN reviews amazon MSE accuracy, interpretability
ISARS (Dang et al., 2021) 2021 CNN, LSTM ratings & reviews amazon RMSE, MAE, NMAE accuracy
MRCP (Liu et al., 2021b) 2021 CNN, Attn ratings & reviews amazon, yelp MSE accuracy
MrRec (Liu et al., 2021d) 2021 CNN, Attn reviews amazon & goodreads MSE accuracy, interpretability
SIFN (Zhang et al., 2021) 2021 Attn reviews amazon MSE accuracy
U-ARM (Sun et al., 2021b) 2021 MLP reviews amazon & yelp RMSE accuracy, explainability
ERP (Wu et al., 2021) 2021 CNN ratings & reviews amazon & yelp MSE, RMSE accuracy
@@
SER (Choi et al., 2022) 2022 MLP reviews amazon & yelp MSE, NDCG accuracy
DRRNN (Xi et al., 2021) 2022 CNN, MLP ratings & reviews amazon RMSE accuracy
KANN (Liu and Miyazaki, 2023) 2022 Attn reviews amazon, Imdb RMSE accuracy, explainability
RI-GCN (Cai et al., 2022) 2022 Graph, CNN ratings & reviews amazon & yelp MSE accuracy
DualGCN (Shi et al., 2022) 2022 Attn reviews amazon & yelp N, MSE, P, R accuracy
RPR (Liu et al., 2022) 2022 CNN reviews amazon & yelp MSE & MAE accuracy, explainability
MAN (Yang et al., 2023c) 2022 CNN, Attn reviews amazon RMSE & MAE accuracy
@@
ExpGCN (Wei et al., 2023) 2023 Graph, CNN ratings & reviews amazon, yelp, trip, Imdb P, R, N accuracy, explainability
MAGCL (Wang et al., 2023b) 2023 Graph & CL ratings & reviews amazon & yelp MRR, NDCG, HR accuracy, interpretability
ARG (Chiang et al., 2023) 2023 RL ratings & reviews amazon & yelp Rel, GF, AE accuracy
LightDL (Chiranjeevi and Rajaram, 2023) 2023 RL reviews amazon & yelp P, R, F accuracy
RMCL (Yang et al., 2023a) 2023 CL reviews amazon MSE accuracy
TSE (Yang et al., 2023b) 2023 CNN reviews amazon MSE, RMSE, MAE accuracy
ERUR (Liu et al., 2023a) 2023 Graph reviews amazon MSE, MAE accuracy
RRAP (Lei et al., 2023) 2023 CNN, LSTM, BERT reviews amazon & yelp MSE accuracy, interpretability
PLLMPAR (Li et al., 2023b) 2023 LLM reviews amazon & yelp MSE, MAE accuracy
DTMRS (Gheewala et al., 2024) 2024 MLP reviews amazon, LibraryThing, yelp MSE, RMSE, MAE accuracy
MCCL (Wei et al., 2024) 2024 GNN & CL reviews amazon & yelp NDCG, P accuracy, interpretability
SERMON (Liao et al., 2024) 2024 CL ratings & reviews amazon, yelp & tripadvisor MSE, BL, RG, BS accuracy, explainability
PRAG (Zhuang et al., 2024) 2024 CL ratings & reviews amazon, yelp & tripadvisor RMSE accuracy, explainability
Note: L = LibraryThings, Eps = Epinions, Rel = Relevance, AE = Attack Effectiveness, BL = BLEU Score, RG = ROUGE, BS = BERTScore
Table 4. The Detailed Summary of the Miscellaneous
Models Year Methodology Features Datasets Evaluation Tasks
AFRR (Xu et al., 2018) 2018 NNF ratings & reviews Imdb & amazon MAP sparsity & transparency
contextCF (Osman et al., 2021) 2021 Similarity ratings & reviews amazon MAE, RMSE, R sparsity & sensitibity
CountER (Tan et al., 2021) 2021 Counterfactual ratings & reviews amazon & yelp NDCG, HR, F1 accuracy & explanation
CRRec (Xiong et al., 2021) 2021 Counterfactual reviews amazon & yelp NDCG, HR, F1 accuracy
LUME (Xv et al., 2022) 2022 Knowledge Dis reviews amazon & yelp MSE, NDCG accuracy & biases
RRRS (Preethi et al., 2017) 2020 Similarity review amazon & yelp MAE accuracy & reliability
TMCSAV (Al-Ghuribi et al., 2023a) 2023 Similarity reviews amazon MSE, RMSE, MAE accuracy & trust
ASCF (Al-Ghuribi et al., 2023b) 2023 K-means clustering reviews amazon & yelp MSE, RMSE, MAE accuracy
Note: NNF = Nearest Neighbour Frequency, Dis = Distillation, Cosine Similarity measure
5.3. Miscellaneous Approaches
This section explores a range of methodologies beyond traditional probabilistic and deep learning approaches. These methods use unique strategies to leverage user reviews, addressing aspects not fully covered in the previously discussed categories. Examples include semantic similarity search, sentiment extraction, counterfactual arguments, reinforcement learning, and nearest neighbor search.
Reinforcement Learning, an agent learns to make decisions by interacting with an environment (Wiering and Van Otterlo, 2012). The agent receives rewards or penalties based on its actions, which guide it toward optimizing its behavior over time. In review-based recommender systems (RBRS), RL can be particularly effective in scenarios where user preferences are dynamic and the system must adapt to continuously evolving data. RL algorithms can optimize the recommendation process by learning from user interactions and feedback, such as reviews, to enhance the personalization and accuracy of the recommendations (Afsar et al., 2022).
@@ 5.3. Miscellaneous Approaches
Counterfactual Explainable Recommendation (CountER) (Tan et al., 2021) leverages counterfactual reasoning to provide clear and meaningful explanations for recommendations by identifying minimal changes in item aspects that could influence user decisions. CountER not only enhances the transparency of recommendations but also assesses the impact and clarity of its explanations, significantly outperforming other methods in explainability.
@@
The Lightweight Unbiased Multi-teacher Ensemble (LUME) (Xv et al., 2022) addresses computational complexity, biased recommendations, and poor generalization by implementing a multi-teacher ensemble with debiased knowledge distillation. This strategy aggregates insights from various pretrained models to form a streamlined, unbiased system that excels in generalization.
On the other hand, some other works focus on utilizing external tools for sentiment score extraction from the reviews, and leveraging them to collaborative filtering method for rating prediction (Musto et al., 2017; Osman et al., 2021; Al-Ghuribi et al., 2020; Huang et al., 2020a). For instance, a multi-criteria recommendation system (Musto et al., 2017) utilizes user reviews to extract aspect-specific sentiment scores through opinion mining, integrating them into collaborative filtering algorithms to enhance prediction accuracy. The Sentiment Aspect-Based Retrieval Engine (SABRE) (Caputo et al., 2017) uses a two-level granularity approach, identifying aspects and sub-aspects within text segments. It performs aspect-based sentiment classification and enables filtering during the exploration of retrieved documents, facilitating fine-grained opinion retrieval. ContextCF (Osman et al., 2021) uses four lexicon-based word lists to extract the context-aware sentiments score from the review, combining them to rating based collaborative filtering for improving the recommendation quality. Additionally, both the Average Trusted Multi-Criteria Similarity (TMCSAV) and Aspect Clustering for Collaborative Filtering (ACCF) (Al-Ghuribi et al., 2020) model employ SEAE (Al-Ghuribi et al., 2021) for semantically enhanced aspect extraction, and generates associated sentiment scores using a lexicon approach, specifically SentiWordNet. While TMCSAV demonstrates that the integration of aspect specific sentiment score improves the performance of the multicriteria recommendation system across all metrics (e.g. MSE, MAE, and RMSE) ACCF is effective in case of unbalanced large-scale reiviews. Similarly, Aspect Sentiment Similarity-based Personalized review Recommendation (A2SPR) (Huang et al., 2020a) generates aspect-specific sentiment score by using SentiWordNet3.0.
In educational settings, a sentiment-based model (Hazar et al., 2022) combines explicit and implicit ratings from user comments to address data sparsity and improve recommendation accuracy. A trusted user model (Al-Ghuribi et al., 2023a), enhances multi-criteria recommendations by incorporating adjective features, reducing prediction errors, and building trust with users. An experimental study (Sumaia et al., 2023) further combines implicit sentiment and aspect information with explicit ratings, demonstrating improved CF performance on large-scale datasets. Additionally, a clustering approach groups similar aspects extracted from reviews to reduce data sparsity, enhancing CF performance in movie recommendations (Al-Ghuribi et al., 2023b). Sentiment analysis also facilitates a distributed recommendation system (Singh et al., 2024) that processes large-scale data for more precise recommendations. In agriculture, a recommendation model (Sharma et al., 2024) applies sentiment analysis and clustering to offer disease-specific product recommendations to farmers, illustrating the versatility of these sentiment-based approaches across different sectors.
The Sentiment Utility Logistic Model (SULM) (Bauman et al., 2017) utilizes sentiment analysis to identify key aspects of products that influence user preferences, enhancing the personalization of recommendations. For example, it might recommend a restaurant specifically for its standout seafood, effectively tailoring suggestions to user preferences. Demonstrated in various settings like restaurants and hotels, SULM’s aspect-focused recommendations have been well-received for improving user satisfaction. Building on this focus on user insights, the Adversarial Training for Review-based Recommendations (ATR) (Rafailidis and Crestani, 2019) model integrates user reviews with collaborative filtering. By using adversarial training, this model aligns deep representations of reviews with user and item features, aiming to improve the accuracy of recommendations. The detailed description of this categories in shown in Table 4.
Context-Specific Suitability of Different Approaches: Each review-based recommender system approach has its own strengths and limitations, making it suitable for specific situations depending on the context and requirements. Probabilistic and topic modeling methods (Blei et al., 2003; Reynolds et al., 2009; Tan et al., 2016) are highly interpretable and effective in handling sparse data. This makes them ideal for scenarios like early-stage products or specialized markets where data is limited but interpretability is crucial. However, they may struggle with capturing complex non-linear relationships in extensive datasets. Deep learning approaches, including CNNs, RNNs, and attention mechanisms (Kim et al., 2016; Zheng et al., 2017; Cheng et al., 2018a; Wu et al., 2021), perform exceptionally well at modeling intricate patterns and non-linear relationships, making them suitable for platforms with abundant data like large e-commerce sites or streaming services. These methods are particularly effective when high accuracy is required, but they often demand significant computational resources and may lack interpretability. Graph Neural Networks (GNNs) (He et al., 2015; Wu et al., 2019c; Liu et al., 2020) are well-suited for applications involving rich relational data, such as social networks, where understanding the interactions between users and items is critical. They can capture high-order connectivity but may be complex to implement. Contrastive learning methods (Khosla et al., 2020; Xia et al., 2022; Yang et al., 2023a) are advantageous in situations with abundant unlabeled data and are effective for cold-start problems, but they require careful tuning and large datasets to perform well. LLM-based approaches (Li et al., 2023b) are powerful for understanding complex language data and capturing detailed semantic relationships in reviews. They are suitable for applications that require advanced natural language understanding but can be resource-intensive and less interpretable. Therefore, the choice of method should align with the application’s specific requirements, such as data availability, the necessity for interpretability, computational constraints, and the complexity of user-item interactions. Table 5 provides a clear summary of the trade-offs between these approaches, highlighting their scalability, interpretability, strengths, limitations, and specific use cases.
Table 5. Comparison of Different Approaches of Review-based Recommender Systems
Approaches Scalability Interp-
retability
Strengths Limitations Specific Use Cases
Probabilistic & Topic Modeling (McAuley and Leskovec, 2013; Tan et al., 2016; Deng et al., 2018) Medium High Effective with sparse data, provides interpretable topics Limited in capturing non-linear relationships; computationally intensive for large datasets E-commerce (e.g., new products), where user history is limited
CNN-Based (Kim et al., 2016; Zheng et al., 2017; Xi et al., 2021) Medium Low-Medium Extracts local features, captures contextual phrases for improved prediction with sparse data Requires substantial data and computational resources, interpretability depends on additional techniques E-commerce sites, especially for products with text-heavy reviews
Attention-Based (Cheng et al., 2018a; Mei et al., 2018; Luo et al., 2021; Lei et al., 2023) Medium High Selectively emphasizes relevant data for high personalization, improves relevance and interpretability May struggle with sparse or superficial reviews, effectiveness relies on review quality Contextual recommendation settings; high-user-interaction platforms
Integrated CNN + Attention (Seo et al., 2017; Chen et al., 2018a; Wang et al., 2021b) Low-Medium Medium-High Combines deep feature extraction with selective focus for better recommendation relevance Complex architectures may increase runtime, challenging to maintain transparency Personalized recommendations with rich review content
RNN-Based (LSTM, GRU) (Da’u and Salim, 2019b; Li and Xu, 2020) Low Medium Processes sequential information for sentiment and aspect progression in reviews Requires detailed sequential data; challenging to interpret, can be computationally demanding Sequential analysis settings; sentiment-based product recommendations
Graph Neural Networks (GNNs) (He et al., 2015; Wu et al., 2019c; Liu et al., 2020) Medium-High Medium-High Captures high-order user-item relationships, handles sparsity and complex interactions Complex data structures; interpretability can be limited without additional techniques Social and item networks; sparse user interactions
Contrastive Learning-Based (Khosla et al., 2020; Xia et al., 2022) Medium Medium Models intricate user-item relationships; effective for multi-aspect connections Limited scalability in large datasets; complex training process Fine-grained preference modeling; cross-modal applications
LLMs-Based (Li et al., 2023b) Low-Medium Low-Medium Extracts specific aspects from textual reviews for enhanced personalization and clarity Data and resource-intensive; interpretability tied to model transparency Rich review platforms; personalized recommendations
Reinforcement Learning-Based (Wiering and Van Otterlo, 2012; Chiang et al., 2023) Low-Medium Low Adapts to dynamic user preferences through continuous feedback, useful in changing environments Requires extensive user interaction data, may struggle with real-time adaptation Dynamic recommendation systems; user-driven content platforms
6. Datasets and Performance of different Models
6.1. Datasets Overview
Review-based recommender models are evaluated mainly with the datasets that contain user, item, and users’ textual review feedback. The detailed statistics and the sources are presented below: